If f(x) = x^2; find f(x+1)
am I on the right track?
f(x+1)^2
(x+1)(x+1)
x^2+2x+1

Answer by Edwin McCravy(8909)

If f(x) = x^2; find f(x+1)
am I on the right track?
f(x+1)^2
(x+1)(x+1)
x^2+2x+1
Yes, that's the correct answer. However you shouldn't 
write the step that I lined through above.  Write it this way:

f(x+1) = 
(x+1)² = 
(x+1)(x+1) = 
x²+2x+1

Edwin